# Characterization of a quantum bus between two driven qubits

### v1

@@ Abstract
We investigate the use of driven qubits coupled to a harmonic oscillator to implement a $\sqrt{i\mathrm{SWAP}}$-gate. By dressing the qubits through an external driving field, the qubits and the harmonic oscillator can be selectively coupled, leading to effective qubit-qubit interactions. We analyze a qubit readout mechanism based on the detection of a shift of the harmonic oscillator’s resonance frequency, and demonstrate that when coupled to low-frequency resonators, dressed qubits provide a more robust readout than bare qubits in the presence of damping and thermal effects. Furthermore, we study the impact of various system parameters on the fidelity of the two qubit-gate, identifying an optimal range for quantum computation. Our findings guide the implementation of high-fidelity quantum gates in experimental setups, for example those employing nanoscale mechanical resonators.
@@ I Introduction
In this work, we investigate the effects of dissipation and thermal effects on dressed qubits coupled to a harmonic oscillator. First, we analyze how the qubit’s state can be inferred by measuring a resonance frequency shift of the harmonic oscillator. In the dispersive regime, such a measurement corresponds to a quantum nondemolition measurement of the qubit. This mechanism has been demonstrated in transmission line resonators [19, 20, 21, 22, 23]. The effect of dissipation on the system due to its coupling to the environment, modeled as a heat bath, depends on the dressing of the qubit. Interestingly, we find that the qualitative form of dissipation differs from that of bare qubits: whereas in bare qubits the stationary state is the ground state, corresponding to one of the qubit eigenstates, a dressed qubit relaxes towards a mixed state. The latter makes the two logical qubit states symmetric in terms of the unwanted dissipation. This may be advantageous since most quantum computing protocols do not differentiate the qualitative properties of the two logical states. We also demonstrate that the state readout mechanism is more robust for a dressed qubit compared to a bare one, particularly at higher temperatures. Following the approach of Refs. [19, 17, 24, 25], we study how the harmonic-oscillator-mediated coupling can be utilized to implement a two-qubit gate. We show how the thermal occupation of the harmonic oscillator affects the ability to generate qubit-qubit entanglement. Additionally, we compare gate fidelities across different model parameters and propose an optimal parameter range for practical implementation. Notably, we show that the fidelity exhibits a non-monotonic dependence on the dressing rise time. This non-monotonic behavior can be leveraged to enhance fidelity in setups with a finite dressing speed.
@@
The work is organized as follows. In Sec. II, we introduce the quantum master equation formalism used to study the qubit-harmonic oscillator system. In Sec. III, we investigate how the damping and temperature affect the qubit state readout mechanism, comparing the performance of dressed and bare qubits and highlighting the advantages of using dressed qubits. In Sec. IV, we analyze how various system parameters influence the fidelity of a two-qubit gate mediated by a harmonic oscillator and propose an optimal parameter range for practical implementation. In Sec. V, we discuss possible realizations of the qubit-harmonic oscillator system across different technological platforms. Finally, we summarize the results in Sec. VI.
@@ II.1 Dressed qubit-harmonic oscillator coupling
We study a qubit with transition frequency $\omega_{q}$, coupled to a harmonic oscillator with resonance frequency $\omega_{h}$. The Hamiltonian describing the coupling is
@@
where $\hat{\sigma}_{i}$ are the Pauli matrices describing the qubit degrees of freedom, $\hat{a}^{\dagger}$ and $\hat{a}$ are the raising and lowering operators of the harmonic oscillator, and $g$ is the coupling strength, which depends on the coupling mechanism and its implementation. As described in Appendix A, within the rotating wave approximation, the coupling term describes exchange of energy quanta between the qubit and the harmonic oscillator. If $\omega_{q}$ is of the order of $\omega_{h}$, the coupling between the subsystems leads to a sizable interaction between them. However, the characteristic frequencies of the qubit and the bus may differ by several orders of magnitude. If the qubit transition frequency $\omega_{q}$ is much higher than $\omega_{h}$, the qubit may be driven by an input field in resonance with $\omega_{q}$ to dress the qubit with the desired eigenfrequency of the Rabi oscillations. For instance, a spin qubit can be dressed by applying a microwave field. Its magnetic field component, pointing in a direction perpendicular to the dc field, dresses the spin qubit, as shown in Fig. 1(a). Here, we treat the driving field as a classical field interacting with the qubit. The Hamiltonian describing the dressed qubit-harmonic oscillator system then reads (here and below $\hbar=k_{B}=1$):
@@
where the first two terms in the Hamiltonian describe the dressed qubit. For a drive in resonance with the bare qubit $\Delta=0$, and within the rotating wave approximation (note that this is distinct from the the rotating wave approximation mentioned for the qubit-harmonic oscillator coupling), Eq. (3) becomes the well-known Jaynes–Cummings Hamiltonian, see Appendix A. The eigenstates of $\hat{\sigma}_{x}$ describe the logical states $\{|0\rangle,|1\rangle\}$ of the qubit. The qubit can be selectively coupled to the harmonic oscillator by changing the driving amplitude $\Omega_{R}(t)$. In Sec. IV.2, we study how the qubit coupling speed affects the fidelity of a harmonic oscillator-mediated gate.
As shown in Appendix A, if the dressed qubit is near-resonant with the harmonic oscillator, $\Omega_{R}\sim\omega_{h}$, the coupling between them shifts the resonance frequency of the oscillator in a way that depends on the qubit state. Therefore, the state of the qubit can be detected by measuring the noise power spectral density of the $\hat{x}(t)=\hat{a}+\hat{a}^{\dagger}$ quadrature
@@ II.2 Quantum master equation
In Eq. (7) we ignore the renormalization of the system Hamiltonian introduced by the system-bath coupling, since the bare parameters are not accessible. $\hat{L}_{k}$ are Lindblad operators describing the effect of the baths on the system. They are related to the system operators coupling to the baths as
@@
where $n_{\mathrm{th},q}=(e^{\omega_{q}/T}-1)^{-1}$ and $n_{\mathrm{th},h}=(e^{\omega_{h}/T}-1)^{-1}$ describe the thermal state of the bath fields coupling to the qubit and the harmonic oscillator, respectively. For the dressed qubit, we assume that the bare transition frequency is much higher than the temperature, i.e., $\omega_{q}\gg T$, so that $n_{\mathrm{th},q}=0$. The terms containing the Lindblad operators describing the coupling of the harmonic oscillator to the heat bath, $\hat{L}_{h,\pm}$, induce damping, driving the harmonic oscillator towards a thermal equilibrium state with the heat bath. Similarly, the terms containing $\hat{L}_{q,\pm}$ thermalize a bare qubit with its environment. Here, we assume that the qubit damping is stronger than pure dephasing, and only consider the Lindblad operators $\hat{L}_{q,\pm}$ describing the damping. In the limit $\omega_{q}\gg T$, the bare qubit evolves towards its ground state in the stationary regime. However, for a dressed qubit, the operators $\hat{\sigma}_{\pm}$ no longer describe energy quanta exchange with the environment, and the system rather evolves towards a mixed state. A dressed qubit ideally decays into a maximally mixed state, but as shown in Sec. III, the coupling with a harmonic oscillator results in a stationary state with a negative $\langle\sigma_{x}\rangle$.
@@
Figure 2: (a) Noise power spectral density for different temperatures for (a) a dressed qubit with $\omega_{q}=50\omega_{h}$, $\Delta=0$ and $\Delta_{R}=0.05\omega_{h}$, and (b) a bare qubit with $\omega_{q}=1.05\omega_{h}$. The solid lines correspond to the initial state $|1\rangle$ and the dashed lines to the initial state $|0\rangle$. The vertical dashed lines indicate the maximum of the spectral density. Time evolution of the qubit state for (c) dressed and (d) bare qubits, for thermal occupation $n_{\mathrm{th},h}=0$. The parameters used for all plots are $g=5\cdot 10^{-3}\omega_{h}$, $\gamma_{q}=10^{-4}\omega_{h}$ and $\gamma_{h}=10^{-4}\omega_{h}$.
III Characterization of the qubit-harmonic oscillator coupling
@@
A quantum circuit consists of a sequence of quantum gates followed by measurements that yield information in the form of a classical bit (0 or 1), according to the Born rule. Qubit measurements are typically made in the qubit’s eigenstate basis, which usually defines the computational basis ${|0\rangle,|1\rangle}$. However, measurements in other bases are also possible by applying appropriate state rotation gates before the measurement. In this section, we present a qubit state readout mechanism that relies on a qubit-dependent shift in the oscillator’s resonance frequency. This approach is similar to how the state of a superconducting qubit is measured by tracking the resonance frequency of a coupled waveguide resonator [19, 30]. We compare the resonance frequency shift for both bare and dressed qubits and analyze how nonzero thermal occupation influences each configuration.
@@
In the dispersive regime $|\Delta_{R}|\gg g$, the resonance frequency of the harmonic oscillator gets shifted by ±$\pm\frac{g^{2}}{\Delta_{R}}$ depending on the qubit’s state [19, 22], see Eq. (A.2). The resonance frequency manifests as a maximum in the noise power spectral density, Eq. (4). The shift of the maximum can be detected, for instance, in an optomechanical quantum bus, through optical measurements, and pulsed measurements allow for the detection of short-lived shifts in the mechanical frequency [28, 31, 32, 16]. The resonance frequency shift is generally measurable if it is greater than the linewidth.
@@
In this section, we analyze the frequency- and time-domain response of a qubit coupled to a harmonic oscillator. We solve the QME (7) for the system Hamiltonian (3) [dressed qubit] and (1) [bare qubit]. The codes used to solve the equation are available at [33]. The temperature $T$ of the heat bath determines the thermal occupations $n_{\mathrm{th},q}$ and $n_{\mathrm{th},h}$ in the Lindblad operators, and the initial state of the harmonic oscillator, which we assume to be in thermal equilibrium with the heat bath so that $\rho_{h}(0)=Z^{-1}e^{-\omega_{h}\hat{a}^{\dagger}\hat{a}/T}$, where $Z$ is the partition function of the harmonic oscillator. The qubit is initialized either in the ground ($|0\rangle$) or the excited ($|1\rangle$) state. The noise power spectral density introduced in Eq. (4) is written in the Heisenberg picture, while the QME formalism is based on the Schrödinger picture. Using the quantum regression theorem [34, 35], the two-time correlation function in Eq. (4) can be rewritten as
@@
In Fig. 2 we show the spectral density of the $\hat{x}$ quadrature [Eq. (4)] for a dressed qubit [Fig. 2(a)] and a bare qubit [Fig. 2(b)]. In order to compute the Fourier transform in Eq. (4), we have considered a finite integration time. Since the system is damped, it eventually reaches a stationary state, which at long integration times dominate the integral for the Fourier transform and thus erases any information from the initial state. This cutoff represents the measurement time in an experiment. We have chosen a cutoff time of $t_{\mathrm{cut}}=40\pi|\Delta_{R}|/g^{2}$, which provides sufficient frequency resolution to accurately determine the position of the peak in the spectral density.
@@
As shown in Fig. 2(a), in the absence of thermal occupation of the oscillator (blue line), the resonance frequency shift for the dressed qubit is given approximately by ±$\pm\frac{g^{2}}{\Delta_{R}}$ [Eq. (A.2)], with the upper and lower signs corresponding to the excited and ground states, respectively. The resonance frequency shift is slightly suppressed at higher temperatures, but the resonance peak height increases, making it easier to detect the resonance frequency. $S(\omega)$ shows a small kink at the frequency corresponding to the opposite state $\mp\frac{g^{2}}{\Delta_{R}}$. This feature can be explained by studying the time-domain response. In panel (c) we show the time evolution of a qubit with initial state $|1\rangle$. The system begins at the excited state with $\langle\sigma_{x}\rangle=1$ (see Hamiltonian (3)), and evolves into a mixed state due to the damping. In the stationary state, the weight of the ground state is higher than that of the excited state, so that $\langle\sigma_{x}\rangle<0$. The contribution from the $|0\rangle$ leads to the kinks of the noise power spectral density at $\omega-\omega_{h}=-\frac{g^{2}}{\Delta_{R}}$.
@@
In panels (b) and (d), we consider a bare qubit with a detuning equivalent to that of the dressed qubit $\omega_{q}=1.05\omega_{h}$. The logical states of the bare qubit are given by the eigenstates of $\hat{\sigma}_{z}$ [see Hamiltonian (1)]. Unlike the dressed qubit, at high temperatures the position of the resonance frequency peak gets sizably shifted to lower values, going below the bare resonance frequency of the oscillator for $n_{\mathrm{th},h}\sim 4$ (yellow line). Moreover, the kink at the opposite frequency becomes very relevant at nonzero thermal occupation, so that the initial state of the qubit cannot be determined from the noise power spectral density for $n_{\mathrm{th},h}\gtrsim 4$. As shown in panel (d), a qubit with initial state $|1\rangle$ rapidly decays to the ground state ($\langle\sigma_{z}\rangle=-1$), leading to a sizable peak at $\omega=\omega_{h}-\frac{g^{2}}{\Delta_{R}}$. For the bare qubit, $\omega_{q}$ needs to be resonant with $\omega_{h}$, so the bare qubit’s quality factor $Q_{q}=\omega_{q}/\gamma_{q}$ is smaller than the dressed qubit’s one, where $\omega_{q}\gg\omega_{h}$. Another drawback of using the bare qubit in resonance with the low-frequency harmonic oscillator is that since $\omega_{q}\sim\omega_{h}$, the qubit thermal occupation $n_{\mathrm{th},q}\sim n_{\mathrm{th,m}}$ is not negligible, leading to a stronger damping effect. Therefore, dressing the qubit does not only provide a mechanism to control the qubit-harmonic oscillator detuning, but it also leads to a lower effective temperature for the qubit. In summary, the dressed qubit is more suitable for the readout mechanism based on the spectral density of the oscillator, since $n_{\mathrm{th},q}$, and consequently the effective damping rate $(n_{\mathrm{th},q}+1/2)\gamma_{q}$, are lower for the dressed qubit than in an bare qubit close to resonance with the oscillator.

### v2

# Characterisation of a quantum bus between two driven qubits

@@ Abstract
We investigate the use of driven qubits coupled to a harmonic oscillator to implement a $\sqrt{i\mathrm{SWAP}}$-gate. By dressing the qubits through an external driving field, the qubits and the harmonic oscillator can be selectively coupled, leading to effective qubit-qubit interactions. We analyze a qubit readout mechanism based on the detection of a shift of the harmonic oscillator’s resonance frequency, and demonstrate that when coupled to low-frequency resonators, dressed qubits provide a more robust readout than bare qubits in the presence of damping and thermal effects. Furthermore, we study the impact of various system parameters on the fidelity of the two-qubit gate, identifying an optimal range for quantum computation. Our findings guide the implementation of high-fidelity quantum gates in experimental setups, for example those employing nanoscale mechanical resonators.
@@ I Introduction
In this work, we investigate the effects of dissipation and thermal effects on dressed qubits coupled to a harmonic oscillator. First, we analyze how the qubit’s state can be inferred by measuring a resonance frequency shift of the harmonic oscillator. In the dispersive regime, such a measurement corresponds to a quantum nondemolition measurement of the qubit. This mechanism has been demonstrated in transmission line resonators [19, 20, 21, 22, 23]. The effect of dissipation on the system due to its coupling to the environment, modeled as a heat bath, depends on the dressing of the qubit. Interestingly, we find that the qualitative form of dissipation differs from that of bare qubits: whereas in bare qubits the steady state is the ground state, corresponding to one of the qubit eigenstates, a dressed qubit relaxes towards a mixed state. The latter makes the two logical qubit states symmetric in terms of the unwanted dissipation. This may be advantageous since most quantum computing protocols do not differentiate the qualitative properties of the two logical states. We also demonstrate that the state readout mechanism is more robust for a dressed qubit compared to a bare one, particularly at higher temperatures. Following the approach of Refs. [19, 17, 24, 25], we study how the harmonic-oscillator-mediated coupling can be utilised to implement a two-qubit gate. We show how the thermal occupation of the harmonic oscillator affects the ability to generate qubit-qubit entanglement. Additionally, we compare gate fidelities across different model parameters and propose an optimal parameter range for practical implementation. Notably, we show that the fidelity exhibits a non-monotonic dependence on the dressing rise time. This non-monotonic behaviour can be leveraged to enhance fidelity in setups with a finite dressing speed.
@@
The work is organised as follows. In Sec. II, we introduce the quantum master equation formalism used to study the qubit-harmonic oscillator system. In Sec. III, we investigate how the damping and temperature affect the qubit state readout mechanism, comparing the performance of dressed and bare qubits and highlighting the advantages of using dressed qubits. In Sec. IV, we analyze how various system parameters influence the fidelity of a two-qubit gate mediated by a harmonic oscillator and propose an optimal parameter range for practical implementation. In Sec. V, we discuss possible realisations of the qubit-harmonic oscillator system across different technological platforms. Finally, we summarise the results in Sec. VI.
@@ II.1 Dressed qubit-harmonic oscillator coupling
We study a qubit with transition frequency $\omega_{q}$, coupled to a harmonic oscillator with resonance frequency $\omega_{h}$. The Hamiltonian describing the coupling is (here and below $\hbar=k_{B}=1$)
@@
where $\hat{\sigma}_{i}$ are the Pauli matrices describing the qubit degrees of freedom, $\hat{a}^{\dagger}$ and $\hat{a}$ are the raising and lowering operators of the harmonic oscillator, and $g$ is the coupling strength, which depends on the coupling mechanism and its implementation. As described in Supplementary note A [29], within the rotating wave approximation, the coupling term describes the exchange of energy quanta between the qubit and the harmonic oscillator. If $\omega_{q}$ is of the order of $\omega_{h}$, the coupling between the subsystems leads to a sizable interaction between them. However, the characteristic frequencies of the qubit and the bus may differ by several orders of magnitude. If the qubit transition frequency $\omega_{q}$ is much higher than $\omega_{h}$, the qubit may be driven by an input field in resonance with $\omega_{q}$ to dress the qubit with the desired eigenfrequency of the Rabi oscillations. For instance, a spin qubit can be dressed by applying a microwave field. Its magnetic field component, pointing in a direction perpendicular to the dc field, dresses the spin qubit, as shown in Fig. 1(a). Here, we treat the driving field as a classical field interacting with the qubit. The Hamiltonian describing the dressed qubit-harmonic oscillator system then reads:
@@
where the first two terms in the Hamiltonian describe the dressed qubit. For a drive in resonance with the bare qubit $\Delta=0$, and within the rotating wave approximation (note that this is distinct from the the rotating wave approximation mentioned for the qubit-harmonic oscillator coupling), Eq. (3) becomes the well-known Jaynes–Cummings Hamiltonian, see Supplementary note A [29]. The eigenstates of $\hat{\sigma}_{x}$ describe the logical states $\{|0\rangle,|1\rangle\}$ of the qubit. The qubit can be selectively coupled to the harmonic oscillator by changing the driving amplitude $\Omega_{R}(t)$. In Sec. IV.2, we study how the qubit coupling speed affects the fidelity of a harmonic oscillator-mediated gate.
As shown in Supplementary note A [29], if the dressed qubit is near-resonant with the harmonic oscillator, $\Omega_{R}\sim\omega_{h}$, the coupling between them shifts the resonance frequency of the oscillator in a way that depends on the qubit state. Therefore, the state of the qubit can be detected by measuring the noise power spectral density of the $\hat{x}(t)=\hat{a}+\hat{a}^{\dagger}$ quadrature
@@ II.2 Quantum master equation
In Eq. (7) we ignore the renormalisation of the system Hamiltonian introduced by the system-bath coupling, since the bare parameters are not accessible. $\hat{L}_{k}$ are Lindblad operators describing the effect of the baths on the system. They are related to the system operators coupling to the baths as
@@
where $n_{\mathrm{th},q}=(e^{\omega_{q}/T}-1)^{-1}$ and $n_{\mathrm{th},h}=(e^{\omega_{h}/T}-1)^{-1}$ describe the thermal state of the bath fields coupling to the qubit and the harmonic oscillator, respectively. For the dressed qubit, we assume that the bare transition frequency is much higher than the temperature, i.e., $\omega_{q}\gg T$, so that $n_{\mathrm{th},q}=0$. The terms containing the Lindblad operators describing the coupling of the harmonic oscillator to the heat bath, $\hat{L}_{h,\pm}$, induce damping, driving the harmonic oscillator towards a thermal equilibrium state with the heat bath. Similarly, the terms containing $\hat{L}_{q,\pm}$ thermalise a bare qubit with its environment. Here, we assume that the qubit damping is stronger than pure dephasing, and only consider the Lindblad operators $\hat{L}_{q,\pm}$ describing the damping.
In the low temperature limit $T\ll\omega_{q}$, the bare qubit evolves towards its ground state in the steady-state regime. However, for a dressed qubit, the operators $\hat{\sigma}_{\pm}$ no longer describe energy quanta exchange with the environment, and the system rather evolves towards a mixed state [31]. A dressed qubit ideally decays into a maximally mixed state, but as shown in Sec. III, the coupling with a harmonic oscillator results in a steady state with a negative $\langle\sigma_{x}\rangle$. The interaction between the qubit and the harmonic oscillator couples the $|1,n=0\rangle$ and $|0,n=1\rangle$ states, where the second index denotes the oscillator’s quantum number. This coupling transfers population from the $|1,n=0\rangle$ state to the $|0,n=1\rangle$ state, enforcing the negative $\langle\sigma_{x}\rangle$.
@@
Figure 2: (a) Noise power spectral density for different temperatures for (a) a dressed qubit with $\omega_{q}=50\omega_{h}$, $\Delta=0$ and $\Delta_{R}=0.05\omega_{h}$, and (b) a bare qubit with $\omega_{q}=1.05\omega_{h}$. The solid lines correspond to the initial state $|1\rangle$ and the dashed lines to the initial state $|0\rangle$. The vertical dashed lines indicate the maximum of the spectral density. Time evolution of the qubit state for (c) dressed and (d) bare qubits, for thermal occupation $n_{\mathrm{th},h}=0$. All plots use the parameters $g=5\cdot 10^{-3}\omega_{h}$, $\gamma_{q}=10^{-4}\omega_{h}$ and $\gamma_{h}=10^{-4}\omega_{h}$.
III Characterisation of the qubit-harmonic oscillator coupling
@@
A quantum circuit consists of a sequence of quantum gates followed by measurements that yield information in the form of a classical bit (0 or 1), according to the Born rule. Qubit measurements are typically made in the qubit’s eigenstate basis, which usually defines the computational basis $\{|0\rangle,|1\rangle\}$. However, measurements in other bases are also possible by applying appropriate state rotation gates before the measurement. In this section, we present a qubit state readout mechanism that relies on a qubit-dependent shift in the oscillator’s resonance frequency. This approach is similar to how the state of a superconducting qubit is measured by tracking the resonance frequency of a coupled waveguide resonator [19, 32]. We compare the resonance frequency shift for both bare and dressed qubits and analyze how nonzero thermal occupation influences each configuration.
@@
In the dispersive regime $|\Delta_{R}|\gg g$, the resonance frequency of the harmonic oscillator gets shifted by ±$\pm\frac{g^{2}}{\Delta_{R}}$ depending on the qubit’s state [19, 22], see Eq. (A.2) in the Supplementary Information [29]. The resonance frequency manifests as a maximum in the noise power spectral density, Eq. (4). The shift of the maximum can be detected, for instance, in an optomechanical quantum bus, through optical measurements, and pulsed measurements allow for the detection of short-lived shifts in the mechanical frequency [28, 33, 34, 16]. The resonance frequency shift is generally measurable if it is greater than the linewidth.
@@
In this section, we analyze the frequency- and time-domain response of a qubit coupled to a harmonic oscillator. We solve the QME (7) for the system Hamiltonian (3) [dressed qubit] and (1) [bare qubit]. The codes used to solve the equation are available at [35]. The temperature $T$ of the heat bath determines the thermal occupations $n_{\mathrm{th},q}$ and $n_{\mathrm{th},h}$ in the Lindblad operators, and the initial state of the harmonic oscillator, which we assume to be in thermal equilibrium with the heat bath so that $\rho_{h}(0)=Z^{-1}e^{-\omega_{h}\hat{a}^{\dagger}\hat{a}/T}$, where $Z$ is the partition function of the harmonic oscillator. The qubit is initialised either in the ground ($|0\rangle$) or the excited ($|1\rangle$) state. The noise power spectral density introduced in Eq. (4) is written in the Heisenberg picture, while the QME formalism is based on the Schrödinger picture. Using the quantum regression theorem [36, 37], the two-time correlation function in Eq. (4) can be rewritten as
@@
In Fig. 2 we show the spectral density of the $\hat{x}$ quadrature [Eq. (4)] for a dressed qubit [Fig. 2(a)] and a bare qubit [Fig. 2(b)]. In order to compute the Fourier transform in Eq. (4), we have considered a finite integration time. Since the system is damped, it eventually reaches a steady state, which at long integration times dominates the integral for the Fourier transform and thus erases any information from the initial state. This cutoff represents the measurement time in an experiment. We have chosen a cutoff time of $t_{\mathrm{cut}}=40\pi|\Delta_{R}|/g^{2}$, which provides sufficient frequency resolution to accurately determine the position of the peak in the spectral density.
@@
As shown in Fig. 2(a), in the absence of thermal occupation of the oscillator (blue line), the resonance frequency shift for the dressed qubit is given approximately by ±$\pm\frac{g^{2}}{\Delta_{R}}$ [Eq. (A.2) in the Supplementary Information [29]], with the upper and lower signs corresponding to the excited and ground states, respectively. The resonance frequency shift is slightly suppressed at higher temperatures, but the resonance peak height increases, making it easier to detect the resonance frequency. $S(\omega)$ shows a small kink at the frequency corresponding to the opposite state $\mp\frac{g^{2}}{\Delta_{R}}$. This feature can be explained by studying the time-domain response. In Fig. 2(c) we show the time evolution of a qubit with initial state $|1\rangle$. The system begins at the excited state with $\langle\sigma_{x}\rangle=1$ (see Hamiltonian (3)), and evolves into a mixed state due to the damping [31]. As shown in Supplementary note B [29], in the steady state, the weight of the ground state is higher than that of the excited state, so that $\langle\sigma_{x}\rangle<0$. The dashed line in Fig. 2(c) indicates the value of $\langle\sigma_{x}\rangle$ in the steady state, given by Eq. (B.4) in the Supplementary Information [29]. The contribution of the $|0\rangle$ state to the noise power spectral density leads to the kinks at $\omega=\omega_{h}-\frac{g^{2}}{\Delta_{R}}$.
@@
In Figs. 2(b) and (d), we consider a bare qubit with a detuning equivalent to that of the dressed qubit $\omega_{q}=1.05\omega_{h}$. The logical states of the bare qubit are given by the eigenstates of $\hat{\sigma}_{z}$ [see Hamiltonian (1)]. Unlike the dressed qubit, at high temperatures the position of the resonance frequency peak gets sizably shifted to lower values, going below the bare resonance frequency of the oscillator for $n_{\mathrm{th},h}\sim 4$ (yellow line). Moreover, the kink at the opposite frequency becomes very relevant at nonzero thermal occupation, so that the initial state of the qubit cannot be determined from the noise power spectral density for $n_{\mathrm{th},h}\gtrsim 4$. As shown in Fig. 2(d), a qubit with initial state $|1\rangle$ rapidly decays to the ground state ($\langle\sigma_{z}\rangle=-1$), leading to a sizable peak at $\omega=\omega_{h}-\frac{g^{2}}{\Delta_{R}}$. For the bare qubit, $\omega_{q}$ needs to be resonant with $\omega_{h}$, so the bare qubit’s quality factor $Q_{q}=\omega_{q}/\gamma_{q}$ is smaller than the dressed qubit’s one, where $\omega_{q}\gg\omega_{h}$. Another drawback of using the bare qubit in resonance with the low-frequency harmonic oscillator is that since $\omega_{q}\sim\omega_{h}$, the qubit thermal occupation $n_{\mathrm{th},q}\sim n_{\mathrm{th,m}}$ is not negligible, leading to a stronger damping effect. Therefore, dressing the qubit does not only provide a mechanism to control the qubit-harmonic oscillator detuning, but it also leads to a lower effective temperature for the qubit. In summary, the dressed qubit is more suitable for the readout mechanism based on the spectral density of the oscillator, since $n_{\mathrm{th},q}$, and consequently the effective damping rate $(n_{\mathrm{th},q}+1/2)\gamma_{q}$, are lower for the dressed qubit than in a bare qubit close to resonance with the oscillator.